# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v2

@@ 1. Patterns where a value is first assigned
Appendix E More Results
@@
Table 16: Cross-Idiom Generalization on Python Ruff Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms, and the performance is reported on other Ruff idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
E.1 Expanded Results on the PEP Hard Idiom Benchmark
@@
We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models. This suggests that SFT alone may lead to overfitting on the Ruff idiom distribution and struggles to generalize from easy to hard cases without DPO training. These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline. However, we also emphasize that while the SFT stage can limit generalization, it remains essential for effective DPO training, as it teaches the LLM to follow the correct output format and establishes a strong base policy. This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
@@
Table 18: Cross-Idiom Generalization on Python Ruff Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on other Ruff idioms with varying levels of transfer (section 4.2). Best score across the compared training setups per model are bolded.
@@
Table 19: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
E.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
@@
Table 25: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings. Each cell shows the test statistic (p-value). All Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
@@
E.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model

### v3

@@ 1. Patterns where a value is first assigned
Appendix G More Results
Table 18: Cross-Idiom Generalization on Python Ruff Best Practice Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms, and the performance is reported on other Ruff best practice idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
@@ Appendix G More Results
Table 19: Ablation on In-Context Examples for the Hard PEP Benchmark: Effect of including before/after code examples in the prompt for base Qwen3-4B models and final MetaLint models trained with SFT and DPO. Across most settings, removing examples results in minimal changes in detection and localization performance, indicating that the natural-language descriptions alone are often sufficient. The only notable degradation occurs for the CoT-trained MetaLint model, which exhibits a modest drop in detection performance without examples due to an increased tendency to predict NO VIOLATIONS, suggesting that this more conservative variant benefits more from additional in-context guidance.
Model Examples? Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B Yes 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B w CoT Yes 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B (SFT + RS-DPO) Yes 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B w CoT (RS-SFT + RS-DPO) Yes 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Qwen3-4B No 0.5556 0.1736 0.2645 0.0802 0.0733 0.0766
Qwen3-4B w CoT No 0.8304 0.3669 0.5089 0.2572 0.1542 0.1928
Qwen3-4B (SFT + RS-DPO) No 0.7167 0.7268 0.7217 0.3396 0.1909 0.2444
Qwen3-4B w CoT (RS-SFT + RS-DPO) No 0.8215 0.4249 0.5601 0.2860 0.1983 0.2342
G.1 Expanded Results on the PEP Hard Best Practice Benchmark
@@
We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models. This suggests that SFT alone may lead to overfitting on the Ruff best practice idiom distribution and struggles to generalize from easy to hard cases without DPO training. These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline. However, we also emphasize that while the SFT stage can limit generalization, it remains essential for effective DPO training, as it teaches the LLM to follow the correct output format and establishes a strong base policy. This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
@@
Table 21: Cross-Idiom Generalization on Python Ruff Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms and tested on other Ruff best practice idioms with varying levels of transfer (section 4.2). Best score across the compared training setups per model is bolded.
@@
Table 22: Easy-to-Hard Generalization on PEP Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms and tested on hard manually curated PEP best practice violation detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
@@
G.2 Statistical Significance of Results on the PEP Hard Best Practice Benchmark
@@
Table 28: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings. Each cell shows the test statistic (p-value). All Qwen2.5 and Qwen2.5Coder variants are instruction-tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
@@
G.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
@@
Table 31: Cross-Idiom Generalization on JEP & PMD Best Practice Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 31). The transfer column indicates training and test data on the left and right side of the arrow. Best score across the compared training setups per model are bolded.
@@
Table 32: Egregious Localization Error Types: Breakdown of recurring error categories observed in a manual analysis of 50 egregious localization failures made by the Qwen3-4B non-CoT model trained with SFT+RS-DPO. An egregious localization error is defined as predicting a non-empty set of line numbers that is completely disjoint from the non-empty ground-truth set, indicating correct detection but failed localization.
Error Type Definition Ocurrences (%)
Adjacent Predicts a line immediately adjacent to the true violating line, but not the correct one. 12%
Before/After Flags a line that follows the best practice rather than the line that violates it. 4%
Definition vs. Initialization Flags an initialization instead of the definition that actually constitutes the violation. 4%
Plausible-Looking Flags a superficially plausible but ultimately incorrect line due to surface-level similarity or insufficient semantic nuance. 46%
Random Selects an apparently arbitrary line, suggesting a collapse in localization despite correct detection. 28%
Repetition Repeats the same incorrect line multiple times, indicating a decoding or sampling artifact. 6%